Mail Index
- [IGSSTATION-3744]: WSLR: new receiver; data publishing resumed
- Date: Thu, 13 May 2010 18:34:00 -0400
- [IGSSTATION-3743]: SEY1: data outage
- Date: Thu, 13 May 2010 16:08:11 -0600
- [IGSSTATION-3741]: KOKB Hydrogen Maser change
- From: Khachikyan, Razmik (3350-Affiliate)
- Date: Thu, 13 May 2010 09:11:11 -0700
- [IGSSTATION-3742]: GA FTP server offline [SEC=UNCLASSIFIED]
- Date: Thu, 13 May 2010 14:49:31 +1000
- [IGSSTATION-3740]: NKLG station status : change of TRIMBLE receiver and antenna
- From: Cardaliaguet Jean-Paul
- Date: Wed, 12 May 2010 20:01:14 +0200
- [IGSSTATION-3739]: SOFI Maybe jamming caused by a UHV radio ?
- Date: Tue, 11 May 2010 13:07:21 +0200
- [IGSSTATION-3738]: MBAR: older data now published
- Date: Mon, 10 May 2010 14:12:09 -0600
- [IGSSTATION-3737]: NTUS: Firmware upgrade LEICA GRX1200PRO
- From: Hans van der Marel - LR
- Date: Mon, 10 May 2010 20:47:11 +0200
- [IGSSTATION-3736]: KUNM: data outage
- Date: Mon, 10 May 2010 08:08:04 -0600
- [IGSSTATION-3735]: GLPS: data publication resuming
- Date: Mon, 10 May 2010 07:58:36 -0600
- [IGSSTATION-3734]: UNBJ, UNB3, UNBN, UNBT, UNBE, and UNBD: Power Outage to Affect UNB GNSS Receivers
- Date: Fri, 7 May 2010 08:53:52 -0300
- [IGSSTATION-3733]: XMIS: ME Upgrade
- Date: Fri, 7 May 2010 11:51:07 +1000
- [IGSSTATION-3732]: QUIN - Radome gone
- Date: Thu, 06 May 2010 19:17:13 -0600
- [IGSSTATION-3731]: BARH new measurement engine 3.019
- Date: Thu, 06 May 2010 09:47:37 -0400
- [IGSSTATION-3730]: GUAM: data outage
- Date: Thu, 06 May 2010 07:27:54 -0600
- [IGSSTATION-3729]: GLPS: data outage
- Date: Thu, 06 May 2010 07:22:44 -0600
- [IGSSTATION-3727]: DARW: ME Upgrade
- Date: Thu, 6 May 2010 15:11:40 +1000
- [IGSSTATION-3728]: ALIC: ME Upgrade
- Date: Thu, 6 May 2010 15:09:55 +1000
- [IGSSTATION-3726]: SALU station is offline
- Date: Wed, 5 May 2010 13:16:53 -0300
- [IGSSTATION-3725]: NSSP: data publication resuming
- Date: Wed, 05 May 2010 08:10:27 -0600
- [IGSSTATION-3724]: KIR0, MAR6, ONSA, SPT0 and VIS0: DOY 124 resubmitted
- Date: Wed, 5 May 2010 15:54:47 +0200
- [IGSSTATION-3723]: HNPT new measurement engine
- Date: Wed, 05 May 2010 08:48:55 -0400
- [IGSSTATION-3722]: MATE Glonass tracking
- From: GeoDAF Administrator
- Date: Wed, 5 May 2010 11:04:10 +0200
- [IGSSTATION-3721]: LAMA: Firmware upgrade
- Date: Wed, 05 May 2010 10:30:37 +0200
- [IGSSTATION-3720]: TOW2: Firmware Upgrade
- Date: Wed, 5 May 2010 13:38:05 +1000
- [IGSSTATION-3719]: STR1: Firmware Upgrade
- Date: Wed, 5 May 2010 10:36:24 +1000
- [IGSSTATION-3718]: SUNM: Decommissioned
- Date: Wed, 5 May 2010 10:27:56 +1000
- [IGSSTATION-3717]: WES2: New measurement engine 3.019 and change in elevation cutoff to 0
- Date: Tue, 04 May 2010 12:31:18 -0400
- [IGSSTATION-3716]: HERT: ME upgrade
- Date: Tue, 4 May 2010 09:01:33 +0100
- [IGSSTATION-3715]: HOB2: Firmware Upgrade
- Date: Tue, 4 May 2010 10:25:46 +1000
- [IGSSTATION-3714]: MAW1: Firmware Upgrade
- Date: Tue, 4 May 2010 10:24:51 +1000
- [IGSSTATION-3713]: CAS1: Firmware Upgrade
- Date: Tue, 4 May 2010 10:21:02 +1000
- [IGSSTATION-3712]: DAV1: Firmware Upgrade
- Date: Tue, 4 May 2010 10:19:09 +1000
- [IGSSTATION-3711]: EPRT updated measurement engine
- Date: Mon, 03 May 2010 17:20:57 -0400
- [IGSSTATION-3710]: SOFI Firmware Upgrade from 7.80/3.018 to 7.80/3.019
- Date: Mon, 03 May 2010 12:08:48 +0200
- [IGSSTATION-3709]: BUCU Firmware Upgrade from 7.80/3.018 to 7.80/3.019
- Date: Mon, 03 May 2010 10:56:25 +0200
- [IGSSTATION-3708]: MOBS: Firmware Upgrade
- Date: Mon, 3 May 2010 17:45:15 +1000
- [IGSSTATION-3707]: WROC: LEICA GRX1200PRO Firmware Upgrade
- Date: Sat, 1 May 2010 11:44:43 +0200
- [IGSSTATION-3706]: KOSG: LEICA GRX1200PRO Firmware Upgrade
- From: Hans van der Marel - LR
- Date: Fri, 30 Apr 2010 23:43:11 +0200
- [IGSSTATION-3705]: Re: BARH Processing engine upgraded to 3.018 wrong engine
- Date: Fri, 30 Apr 2010 16:12:36 -0400
- [IGSSTATION-3704]: BARH Processing engine upgraded to 3.018
- Date: Fri, 30 Apr 2010 16:06:44 -0400
- [IGSSTATION-3703]: NANO: GLONASS tracking resumed
- Date: Fri, 30 Apr 2010 14:51:28 -0400
- [IGSSTATION-3702]: MBAR: station repaired
- Date: Fri, 30 Apr 2010 11:17:45 -0600
- [IGSSTATION-3701]: NKLG station status
- From: Cardaliaguet Jean-Paul
- Date: Fri, 30 Apr 2010 17:27:22 +0200
- [IGSSTATION-3700]: KHAR: receiver firmware upgrade
- Date: Fri, 30 Apr 2010 17:45:04 +0300 (EEST)
- [IGSSTATION-3699]: Leica GRX1200 GG Pro Receivers - resolution
- Date: Fri, 30 Apr 2010 16:05:26 +0200
- [IGSSTATION-3698]: BUCU Resubmission of the daily data file DOY 119
- Date: Fri, 30 Apr 2010 15:46:51 +0200
- [IGSSTATION-3697]: SELE: operational again
- Date: Fri, 30 Apr 2010 07:27:51 -0600
- [IGSSTATION-3696]: Old GLONASS-714 reactivated
- Date: Fri, 30 Apr 2010 14:40:14 +0200
- [IGSSTATION-3695]: SOFI Replacement of receiver and antenna
- Date: Fri, 30 Apr 2010 14:23:04 +0200
- [IGSSTATION-3694]: GLSV: receiver firmware upgrade
- Date: Fri, 30 Apr 2010 14:43:28 +0300 (EEST)
- [IGSSTATION-3693]: HERT: GRX1200GGPRO tracking problem
- Date: Fri, 30 Apr 2010 08:02:30 +0100
- [IGSSTATION-3692]: NANO: GPS tracking resumed and firmware upgraded
- Date: Thu, 29 Apr 2010 17:48:21 -0400
- [IGSSTATION-3691]: BRFT new firmware and changed elevation cutoff angle
- Date: Thu, 29 Apr 2010 15:51:22 -0400
- [IGSSTATION-3690]: FLIN: data outage
- Date: Thu, 29 Apr 2010 14:24:05 -0400
- [IGSSTATION-3689]: NANO: loss of tracking
- Date: Thu, 29 Apr 2010 12:25:49 -0400
- [IGSSTATION-3688]: WES2 EPRT GLONASS tracking switched off
- Date: Thu, 29 Apr 2010 08:55:25 -0400
- [IGSSTATION-3687]: Tracking Problems
- Date: Thu, 29 Apr 2010 22:43:40 +1000
- [IGSSTATION-3686]: GLSV: Firmware upgrade
- Date: Thu, 29 Apr 2010 14:41:57 +0300 (EEST)
- [IGSSTATION-3685]: MATE, VEN1 and USAL
- From: GeoDAF Administrator
- Date: Thu, 29 Apr 2010 11:02:45 +0200
- [IGSSTATION-3684]: Leica GRX1200 GG Pro Receivers
- Date: Thu, 29 Apr 2010 10:55:56 +0200
- [IGSSTATION-3683]: MATE, VEN1 and USAL permanent stations (Leica receiver)
- From: GeoDAF Administrator
- Date: Thu, 29 Apr 2010 10:08:03 +0200
- [IGSSTATION-3682]: Tracking Problems
- Date: Thu, 29 Apr 2010 17:39:26 +1000
- [IGSSTATION-3681]: WROC: Receiver malfunction, loss of data
- Date: Thu, 29 Apr 2010 09:30:41 +0200
- [IGSSTATION-3680]: HILO Decomissioned
- Date: Wed, 28 Apr 2010 15:31:06 -1000
- [IGSSTATION-3679]: CUSV: recent data outages
- Date: Wed, 28 Apr 2010 09:39:12 -0600
- [IGSSTATION-3678]: SELE: data outage
- Date: Tue, 27 Apr 2010 07:44:14 -0600
- [IGSSTATION-3677]: HOLM data doss Day 113
- Date: Fri, 23 Apr 2010 21:16:57 -0400
- [IGSSTATION-3676]: MKEA: station down
- Date: Thu, 22 Apr 2010 09:06:00 -0600
- [IGSSTATION-3675]: LHAZ Resubmission of the daily data file DOY 110
- Date: Wed, 21 Apr 2010 09:14:58 +0200
- [IGSSTATION-3674]: SEY1: data publication resuming
- Date: Tue, 20 Apr 2010 07:33:18 -0600
- [IGSSTATION-3673]: LHAZ Resubmission of the daily data file DOY 109
- Date: Tue, 20 Apr 2010 14:10:00 +0200
- [IGSSTATION-3672]: CMP9: receiver swap
- Date: Mon, 19 Apr 2010 14:55:58 -0600
- [IGSSTATION-3671]: SOFI Malfunctioning Antenna / Shipment delay due to aircraft grounding in Central Europe
- Date: Mon, 19 Apr 2010 13:07:13 +0200
- [IGSSTATION-3670]: ISTA Submission of missing daily data files DOY 099 and DOY 104, summary of unavailable data
- Date: Mon, 19 Apr 2010 12:34:31 +0200
- [IGSSTATION-3669]: LHAZ Resubmission of the daily data file DOY 097
- Date: Mon, 19 Apr 2010 12:02:49 +0200
- [IGSSTATION-3668]: BUCU Resubmission of the daily data file DOY 100
- Date: Mon, 19 Apr 2010 11:15:24 +0200
- [IGSSTATION-3667]: LHAZ Resubmission of the daily data file DOY 107
- Date: Mon, 19 Apr 2010 11:09:20 +0200
- [IGSSTATION-3666]: THTI station status
- Date: Mon, 19 Apr 2010 09:32:59 +0200
- [IGSSTATION-3665]: ONSA resubmitted, DoY 107
- Date: Sun, 18 Apr 2010 11:15:19 +0200
- [IGSSTATION-3664]: WES2 firmware upgrade 7.5 to 7.8
- Date: Fri, 16 Apr 2010 12:51:17 -0400
- [IGSSTATION-3663]: CHWK: Data postings resume
- Date: Wed, 14 Apr 2010 18:31:47 -0400
- [IGSSTATION-3662]: RBAY: data publication resuming
- Date: Wed, 14 Apr 2010 07:24:06 -0600
- [IGSSTATION-3661]: SEY1: data outage
- Date: Tue, 13 Apr 2010 13:33:27 -0600
- [IGSSTATION-3660]: CMP9: station down
- Date: Tue, 13 Apr 2010 13:30:25 -0600
- [IGSSTATION-3659]: TUBO: data gap 10/099-10/102
- Date: Mon, 12 Apr 2010 20:45:49 +0200 (CEST)
- [IGSSTATION-3658]: RBAY: station down
- Date: Mon, 12 Apr 2010 09:01:30 -0600
- [IGSSTATION-3657]: GOPE: data gap 10/101-10/102
- Date: Mon, 12 Apr 2010 14:09:06 +0200 (CEST)
- [IGSSTATION-3656]: SYOG data outage
- Date: Mon, 12 Apr 2010 15:40:28 +0900
- [IGSSTATION-3655]: ISTA Resubmission of the daily data file DOY 098
- Date: Fri, 09 Apr 2010 15:05:42 +0200
- [IGSSTATION-3654]: MORP data outage
- Date: Thu, 8 Apr 2010 10:16:43 +0100
- [IGSSTATION-3653]: FALK: data publication resuming
- Date: Wed, 07 Apr 2010 13:03:51 -0600
- [IGSSTATION-3652]: KOKB: data outage
- Date: Wed, 07 Apr 2010 07:51:37 -0600
- [IGSSTATION-3651]: SOFI Malfunctioning Receiver
- Date: Wed, 07 Apr 2010 15:36:37 +0200
- [IGSSTATION-3650]: ALBH: data flow resumed
- Date: Tue, 6 Apr 2010 13:35:54 -0400
- [IGSSTATION-3649]: R23 almanac problem fixed
- Date: Tue, 6 Apr 2010 15:04:31 +0200
- [IGSSTATION-3648]: GLPS: data publication resuming
- Date: Mon, 05 Apr 2010 12:50:22 -0600
- [IGSSTATION-3647]: ALBH: data interruption effective 2010-D092
- Date: Sun, 4 Apr 2010 14:27:37 -0400
- [IGSSTATION-3646]: CHWK: Data Outage Commenicing 2010-088
- Date: Thu, 1 Apr 2010 17:18:44 -0400
- [IGSSTATION-3645]: HOFN Resubmission of the daily data file DOY 089
- Date: Thu, 01 Apr 2010 14:32:57 +0200
- [IGSSTATION-3644]: NSSP: data outage
- Date: Wed, 31 Mar 2010 07:51:19 -0600
- [IGSSTATION-3643]: SOFI Firmware Downgrade
- Date: Wed, 31 Mar 2010 13:19:39 +0200
- [IGSSTATION-3642]: HOFN Firmware Downgrade
- Date: Wed, 31 Mar 2010 12:50:40 +0200
- [IGSSTATION-3641]: R23 tracking failure concerning JPS/TPS receiver models
- Date: Wed, 31 Mar 2010 12:21:56 +0200
- [IGSSTATION-3640]: MDO1: station repaired
- Date: Tue, 30 Mar 2010 08:08:22 -0600
- [IGSSTATION-3639]: HERT: Firmware upgrade
- Date: Tue, 30 Mar 2010 14:40:54 +0100
- [IGSSTATION-3638]: HLFX, SCH2, VALD, WHIT, BAKE resubmitted, DoY 086
- Date: Mon, 29 Mar 2010 09:42:54 -0400
- [IGSSTATION-3637]: IENG: hourly and daily files missing
- Date: Fri, 26 Mar 2010 18:23:29 +0100
- [IGSSTATION-3636]: OHI2 Firmware Upgrade
- Date: Fri, 26 Mar 2010 15:54:46 +0100
- [IGSSTATION-3635]: HOFN Firmware Upgrade
- Date: Fri, 26 Mar 2010 14:54:08 +0100
- [IGSSTATION-3634]: LHAZ Firmware Upgrade
- Date: Fri, 26 Mar 2010 14:21:32 +0100
- [IGSSTATION-3633]: REYK Firmware Upgrade
- Date: Fri, 26 Mar 2010 13:51:11 +0100
- [IGSSTATION-3632]: SOFI Firmware Upgrade
- Date: Fri, 26 Mar 2010 13:22:53 +0100
- [IGSSTATION-3631]: WTZZ Firmware Upgrade
- Date: Fri, 26 Mar 2010 12:52:23 +0100
- [IGSSTATION-3629]: many bogus PRN25 nav records still
- Date: Thu, 25 Mar 2010 12:45:25 -0400
- [IGSSTATION-3628]: Hourly data from thu3 and qaq1 now available on OLG
- From: Ole Bjerregaard Hansen
- Date: Thu, 25 Mar 2010 16:05:54 +0100
- [IGSSTATION-3627]: USNO data available shortly
- Date: Thu, 25 Mar 2010 10:05:40 -0400
- [IGSSTATION-3630]: Decommissioning SUNM
- Date: Thu, 25 Mar 2010 14:55:03 +1000
- [IGSSTATION-3626]: SUTH: data publication resuming
- Date: Wed, 24 Mar 2010 08:24:49 -0600
- [IGSSTATION-3625]: TLSE station status : satellite system
- From: Cardaliaguet Jean-Paul
- Date: Tue, 23 Mar 2010 16:52:28 +0100
- [IGSSTATION-3624]: MDO1: data outage
- Date: Tue, 23 Mar 2010 08:14:00 -0600
- [IGSSTATION-3623]: BRFT still offline
- Date: Mon, 22 Mar 2010 17:42:33 -0400
- [IGSSTATION-3622]: SUTH: data outage
- Date: Mon, 22 Mar 2010 13:23:33 -0600
- [IGSSTATION-3621]: EPRT upgraded firmware
- Date: Mon, 22 Mar 2010 13:03:39 -0400
- [IGSSTATION-3620]: ASI Italian LDC was down during the last weekend
- From: GeoDAF Administrator
- Date: Mon, 22 Mar 2010 12:56:55 +0100
- [IGSSTATION-3619]: PARC equipment change
- Date: Sun, 21 Mar 2010 11:29:14 -0500
- [IGSSTATION-3618]: STJO change of equipment
- Date: Thu, 18 Mar 2010 15:22:36 -0400
- [IGSSTATION-3617]: MDO1: data outage
- Date: Thu, 18 Mar 2010 11:48:41 -0600
- [IGSSTATION-3616]: BREW: data collection has resumed
- Date: Thu, 18 Mar 2010 08:04:30 -0600
- [IGSSTATION-3615]: MKEA: data gap
- Date: Thu, 18 Mar 2010 07:45:35 -0600
- [IGSSTATION-3614]: MATE configuration change
- From: GeoDAF Administrator
- Date: Wed, 17 Mar 2010 13:51:03 +0100
- [IGSSTATION-3613]: FALK: data outage
- Date: Tue, 16 Mar 2010 14:12:12 -0600
- [IGSSTATION-3612]: CNES Operational Data Center temporary unavailable
- From: Cardaliaguet Jean-Paul
- Date: Tue, 16 Mar 2010 16:52:26 +0100
- [IGSSTATION-3611]: CNES Operational Data Center temporary unavailable
- From: Cardaliaguet Jean-Paul
- Date: Tue, 16 Mar 2010 16:43:49 +0100
- [IGSSTATION-3610]: RBAY: data publication resuming
- Date: Mon, 15 Mar 2010 14:03:42 -0600
- [IGSSTATION-3609]: NOT1 reference clock
- From: GeoDAF Administrator
- Date: Mon, 15 Mar 2010 11:27:23 +0100
- [IGSSTATION-3608]: BREW: data outage
- Date: Fri, 12 Mar 2010 14:10:10 -0700
- [IGSSTATION-3607]: KIR0, MAR6, VIS0 and ONSA resubmitted, DoY 069
- Date: Fri, 12 Mar 2010 15:58:09 +0100
- [IGSSTATION-3606]: STR1: Clock Change
- Date: Fri, 12 Mar 2010 10:12:19 +1100
- [IGSSTATION-3604]: CONZ Daily and hourly data files DOY 058, 059, 060, 061, 062, 063, 064, 065, 066, 067
- Date: Tue, 09 Mar 2010 15:30:04 +0100
- [IGSSTATION-3605]: ELBA data flow resumed
- From: GeoDAF Administrator
- Date: Tue, 9 Mar 2010 09:14:52 +0100
- [IGSSTATION-3603]: KIR0, MAR6, VIS0 and ONSA resubmitted, DoY 067
- Date: Tue, 9 Mar 2010 08:58:35 +0100
- [IGSSTATION-3602]: new VALP data at SOPAC
- Date: Mon, 8 Mar 2010 12:07:16 -0500 (EST)
- [IGSSTATION-3601]: NOT1 reference clock
- From: GeoDAF Administrator
- Date: Fri, 5 Mar 2010 11:53:18 +0100
- [IGSSTATION-3600]: MEDI reference clock
- From: GeoDAF Administrator
- Date: Fri, 5 Mar 2010 11:47:45 +0100
- [IGSSTATION-3599]: 2010 045 Erroneous data uploaded to IGS distribution centers
- From: Weingarth Leonard A Mr NGA-SNGR USA CIV
- Date: Wed, 3 Mar 2010 17:02:53 -0600
- [IGSSTATION-3598]: YKRO: data publication resuming
- Date: Wed, 03 Mar 2010 14:09:50 -0700
- [IGSSTATION-3597]: VALP ANTC and COPO status
- Date: Tue, 2 Mar 2010 11:13:18 -0500 (EST)
- [IGSSTATION-3596]: GLPS: data outage
- Date: Tue, 02 Mar 2010 08:40:39 -0700
- [IGSSTATION-3595]: MAR6: snow
- Date: Tue, 2 Mar 2010 13:14:46 +0100
- [IGSSTATION-3594]: NRL1 power failure
- Date: Mon, 01 Mar 2010 14:55:11 -0500
- [IGSSTATION-3593]: WTZS The ASH701945G_M antenna replaced with a LEIAR25.R3 antenna
- Date: Mon, 01 Mar 2010 12:14:10 +0100
- [IGSSTATION-3592]: CONT Data files available until DOY 058 19:59 UT after earthquake
- Date: Mon, 01 Mar 2010 10:29:08 +0100
- [IGSSTATION-3591]: CONZ Data files available until DOY 058 19:59 UT after earthquake
- Date: Mon, 01 Mar 2010 10:28:08 +0100
- [IGSSTATION-3590]: ESTI: Decommissioned 2003-057
- Date: Fri, 26 Feb 2010 11:12:47 -0500
- [IGSSTATION-3589]: IENG: UTC(IT) Master Clock (Active Hydrogen Maser) replacement
- Date: Fri, 26 Feb 2010 16:11:43 +0100
- [IGSSTATION-3588]: BRFT not sending data
- Date: Fri, 26 Feb 2010 06:59:21 -0500
- [IGSSTATION-3587]: MANA new log submitted
- Date: Thu, 25 Feb 2010 14:59:53 -0500
- [IGSSTATION-3586]: OPMT: missing GPS data today
- Date: Thu, 25 Feb 2010 17:24:30 +0100
- [IGSSTATION-3585]: THTI station status
- From: Cardaliaguet Jean-Paul
- Date: Thu, 25 Feb 2010 09:00:30 +0100
- [IGSSTATION-3584]: KERG station status
- From: Cardaliaguet Jean-Paul
- Date: Thu, 25 Feb 2010 08:53:23 +0100
- [IGSSTATION-3583]: MANA new receiver
- Date: Wed, 24 Feb 2010 15:27:23 -0500
- [IGSSTATION-3582]: TLSE station status : change antenna
- Date: Wed, 24 Feb 2010 17:05:26 +0100
- [IGSSTATION-3581]: YKRO: data outage
- Date: Wed, 24 Feb 2010 07:59:36 -0700
- [IGSSTATION-3580]: LHAZ Submission of the daily data file DOY 053
- Date: Tue, 23 Feb 2010 14:10:32 +0100
- [IGSSTATION-3579]: CONZ Resubmission of the daily data file DOY 053
- Date: Tue, 23 Feb 2010 13:31:33 +0100
- [IGSSTATION-3578]: SOFI Resubmission of the daily data file DOY 053
- Date: Tue, 23 Feb 2010 13:28:38 +0100
- [IGSSTATION-3577]: MARS - Receiver problem
- Date: Mon, 22 Feb 2010 13:15:11 +0100
- [IGSSTATION-3576]: THTI station status
- From: Cardaliaguet Jean-Paul
- Date: Mon, 22 Feb 2010 12:27:48 +0100
- [IGSSTATION-3575]: MEDI reference clcock
- From: GeoDAF Administrator
- Date: Mon, 22 Feb 2010 11:25:52 +0100
- [IGSSTATION-3573]: MEDI frequency standard
- From: GeoDAF Administrator
- Date: Thu, 18 Feb 2010 16:28:38 +0100
- [IGSSTATION-3574]: Status of MANA station (~1 meter East shift starting 2010 040)
- Date: Thu, 18 Feb 2010 06:13:42 -0500
- [IGSSTATION-3572]: LAMP station up and running
- From: GeoDAF Administrator
- Date: Thu, 18 Feb 2010 11:31:27 +0100
- [IGSSTATION-3571]: WSLR: communication problems
- Date: Wed, 17 Feb 2010 16:11:25 -0500
- [IGSSTATION-3570]: IENG: hourly and daily files missing
- Date: Wed, 17 Feb 2010 11:30:54 +0100
- [IGSSTATION-3569]: LPAL: firmware upgraded
- From: Gonzalo López Pedro
- Date: Wed, 17 Feb 2010 10:10:10 +0100
- [IGSSTATION-3568]: BUCU Resubmission of the daily data file DOY 047
- Date: Wed, 17 Feb 2010 09:52:27 +0100
- [IGSSTATION-3567]: STR1: Cesium Offline
- Date: Wed, 17 Feb 2010 16:07:02 +1100 (EST)
- [IGSSTATION-3566]: POL2: data outage
- Date: Tue, 16 Feb 2010 09:28:18 -0700
- [IGSSTATION-3565]: MEDI frequency standard
- From: GeoDAF Administrator
- Date: Tue, 16 Feb 2010 16:21:25 +0100
- [IGSSTATION-3564]: VALD tracking resumed
- Date: Mon, 15 Feb 2010 14:44:14 -0500
- [IGSSTATION-3563]: DUBO outage
- Date: Fri, 12 Feb 2010 09:49:06 -0500
- [IGSSTATION-3562]: LHAZ Resubmission of the daily data file DOY 042
- Date: Fri, 12 Feb 2010 09:06:21 +0100
- [IGSSTATION-3561]: JPLM: data gap
- Date: Wed, 10 Feb 2010 07:43:45 -0700
- [IGSSTATION-3560]: OHI3 Upgrade of the Firmware from 7.53/3.018 to 7.80/3.018
- Date: Tue, 09 Feb 2010 16:28:05 +0100
- [IGSSTATION-3559]: CONZ Upgrade of the firmware from 7.50/3.018 to 7.80/3.018
- Date: Tue, 09 Feb 2010 15:53:03 +0100
- [IGSSTATION-3558]: WROC station firmware update
- Date: Tue, 9 Feb 2010 15:43:44 +0100
- [IGSSTATION-3557]: NICO Upgrade of the firmware from 7.50/3.018 to 7.80/3.018
- Date: Tue, 09 Feb 2010 15:26:01 +0100
- [IGSSTATION-3556]: BUCU Upgrade of the firmware from 7.50/3.018 to 7.80/3.018
- Date: Tue, 09 Feb 2010 14:57:24 +0100
- [IGSSTATION-3555]: ORID Upgrade of the firmware from 7.50/3.018 to 7.80/3.018
- Date: Tue, 09 Feb 2010 14:02:26 +0100
- [IGSSTATION-3554]: WTZR Upgrade of the firmware from 7.53/3.018 to 7.80/3.108
- Date: Tue, 09 Feb 2010 09:47:11 +0100
- [IGSSTATION-3553]: TUKT data submissions resume
- Date: Tue, 9 Feb 2010 00:23:35 -0500
- [IGSSTATION-3552]: ALRT antenna snow removal 2010-01-28 & 2010-02-08
- Date: Mon, 8 Feb 2010 13:37:35 -0500
- [IGSSTATION-3551]: BRST -
- Date: Mon, 08 Feb 2010 14:15:23 +0100
- [IGSSTATION-3550]: CONZ Conformity of the Header in the meteorological data files and the Site Information Form (sitelog)
- Date: Mon, 08 Feb 2010 13:21:02 +0100
- [IGSSTATION-3549]: LHAZ Resubmission of the daily data file DOY 036
- Date: Mon, 08 Feb 2010 12:05:42 +0100
- [IGSSTATION-3548]: BRST - temporarly down
- Date: Mon, 08 Feb 2010 09:47:04 +0100
- [IGSSTATION-3547]: CAGS data loss Days 035-037
- Date: Sun, 7 Feb 2010 13:21:05 -0500
- [IGSSTATION-3546]: DRAO: External Clock Change Completed
- Date: Sat, 6 Feb 2010 14:48:46 -0500
- [IGSSTATION-3545]: BILL: site log correction
- Date: Fri, 05 Feb 2010 11:16:49 -0700
- [IGSSTATION-3544]: GLPS: station repaired
- Date: Thu, 04 Feb 2010 15:28:24 -0700
- [IGSSTATION-3543]: CLAR: data outage correction
- Date: Thu, 04 Feb 2010 10:51:40 -0700
- [IGSSTATION-3542]: SEY1: data publication resuming
- Date: Thu, 04 Feb 2010 08:45:08 -0700
- [IGSSTATION-3541]: CLAR: receiver swap
- Date: Thu, 04 Feb 2010 08:35:37 -0700
- [IGSSTATION-3540]: DRAO: Clock Change(s)
- Date: Wed, 3 Feb 2010 19:03:10 -0500
- [IGSSTATION-3539]: WES2: Data flowing again!
- Date: Wed, 03 Feb 2010 18:14:05 -0500
- [IGSSTATION-3538]: GLPS: station down
- Date: Wed, 03 Feb 2010 15:50:04 -0700
- [IGSSTATION-3537]: IENG: measurements generation resumed
- Date: Tue, 2 Feb 2010 15:18:06 +0100
- [IGSSTATION-3536]: CONZ Firmware Upgrade from 7.50/3.017 to 7.50/3.018 and change in the antenna serial number
- Date: Tue, 02 Feb 2010 14:47:14 +0100
- [IGSSTATION-3535]: PTBB External frequency changed to active H-Maser
- Date: Tue, 02 Feb 2010 10:52:29 +0100
- [IGSSTATION-3534]: YKRO: data publication resuming
- Date: Mon, 01 Feb 2010 14:49:21 -0700
- [IGSSTATION-3533]: OHI3 Upgrade of the firmware from 7.53/3.017 to 7.53/3.018
- Date: Mon, 01 Feb 2010 17:17:45 +0100
- [IGSSTATION-3532]: NICO Upgrade of the firmware from 7.50/3.017 to 7.50/3.018
- Date: Mon, 01 Feb 2010 15:29:48 +0100
- [IGSSTATION-3531]: ORID Upgrade of the firmware from 7.50/3.017 to 7.50/3.018
- Date: Mon, 01 Feb 2010 14:59:16 +0100
- [IGSSTATION-3530]: BUCU Upgrade of the firmware from 7.50/3.017 to 7.50/3.018
- Date: Mon, 01 Feb 2010 14:09:37 +0100
- [IGSSTATION-3529]: SCH2: position shift and snow removal
- Date: Fri, 29 Jan 2010 19:52:47 -0500
- [IGSSTATION-3528]: TUKT communications down since Day 028
- Date: Fri, 29 Jan 2010 17:58:07 -0500
- [IGSSTATION-3527]: WTZR Upgrade of the firmware from 7.53/3.017 to 7.53/3.018
- Date: Fri, 29 Jan 2010 15:55:28 +0100
- [IGSSTATION-3526]: OHI3 Resubmission of the daily data file DOY 026
- Date: Wed, 27 Jan 2010 16:03:16 +0100
- [IGSSTATION-3525]: LBCH: station repaired
- Date: Wed, 27 Jan 2010 07:37:18 -0700
- [IGSSTATION-3524]: TIDB resubmitted
- From: Khachikyan, Razmik (3350-Affiliate)
- Date: Tue, 26 Jan 2010 12:08:53 -0800
- [IGSSTATION-3523]: GLPS: network connection repaired
- Date: Tue, 26 Jan 2010 10:50:20 -0700
- [IGSSTATION-3522]: OHI3 Resubmission of data files DOY 022, 023, 024 and 025
- Date: Tue, 26 Jan 2010 12:54:06 +0100
- [IGSSTATION-3521]: BOGI: receiver change
- Date: Tue, 26 Jan 2010 12:21:48 +0100
- [IGSSTATION-3520]: OHI3 Antenna height for the new LEIAR25.R3 antenna, already installed 2010-01-22 19:45 UT
- Date: Tue, 26 Jan 2010 11:31:04 +0100
- [IGSSTATION-3519]: ALRT antenna snow removal 2010-01-25
- Date: Mon, 25 Jan 2010 11:12:50 -0500
- [IGSSTATION-3518]: OHI3 New Leica AR25 Revision 3 antenna in OHI3 O'Higgins / Antarctica
- Date: Mon, 25 Jan 2010 15:49:48 +0100
- [IGSSTATION-3517]: LBCH: data outage
- Date: Thu, 21 Jan 2010 10:56:38 -0700
- [IGSSTATION-3516]: GLPS: data outage
- Date: Wed, 20 Jan 2010 08:06:45 -0700
- [IGSSTATION-3515]: IISC: recent data gap
- Date: Tue, 19 Jan 2010 16:14:42 -0700
- [IGSSTATION-3514]: YKRO: data outage
- Date: Tue, 19 Jan 2010 16:01:51 -0700
- [IGSSTATION-3513]: BILL: site log update, data publication resuming
- Date: Tue, 19 Jan 2010 15:32:55 -0700
- [IGSSTATION-3512]: DUBO dataflow resumed
- Date: Tue, 19 Jan 2010 11:45:37 -0500
- [IGSSTATION-3511]: SCH2 corrupted data on DOY015
- Date: Mon, 18 Jan 2010 15:44:51 -0500
- [IGSSTATION-3510]: YKRO: data publication resuming
- Date: Thu, 14 Jan 2010 09:51:17 -0700
- [IGSSTATION-3509]: HOLB: Data flow restored
- Date: Thu, 14 Jan 2010 11:33:00 -0500
- [IGSSTATION-3508]: COCO site log update [SEC=UNCLASSIFIED]
- Date: Thu, 14 Jan 2010 11:13:33 +1100 (EST)
- [IGSSTATION-3507]: MOBS receiver problems [SEC=UNCLASSIFIED]
- Date: Thu, 14 Jan 2010 09:28:19 +1100 (EST)
- [IGSSTATION-3506]: NAIN back on-line
- Date: Wed, 13 Jan 2010 16:08:51 -0500
- [IGSSTATION-3505]: ALRT antenna snow removal 2010-01-11
- Date: Tue, 12 Jan 2010 18:21:35 -0500
- [IGSSTATION-3504]: HARV: data outage
- Date: Tue, 12 Jan 2010 08:56:54 -0700
- [IGSSTATION-3503]: ISPA: data outage
- Date: Tue, 12 Jan 2010 08:03:35 -0700
- [IGSSTATION-3502]: RE: [IGSSTATION-3501]: ITRF 2005 coordinates.
- Date: Mon, 11 Jan 2010 15:17:37 -0500
- [IGSSTATION-3499]: MORP: clearance of snow & removal of met sensor
- Date: Mon, 11 Jan 2010 11:53:23 +0000
- [IGSSTATION-3500]: NAIN off-line since 2010-01-06, not 2009-01-06
- Date: Sun, 10 Jan 2010 20:27:31 -0500
- [IGSSTATION-3498]: NAIN off-line since 2009-01-06 (D006)
- Date: Sun, 10 Jan 2010 20:11:47 -0500
- [IGSSTATION-3501]: ITRF 2005 coordinates.
- From: Directeur Technique de l' Instititut Géographique du Burkina
- Date: Sun, 10 Jan 2010 10:30:55 -0000
- [IGSSTATION-3497]: ALGO dataflow resumed
- Date: Fri, 8 Jan 2010 16:07:12 -0500
- [IGSSTATION-3496]: FTP access to RINEX data from JPL will be down next week (Jan 11-17)
- From: Stowers, David A (3350)
- Date: Thu, 7 Jan 2010 10:44:17 -0800
- [IGSSTATION-3495]: ALGO dataflow stop
- Date: Thu, 7 Jan 2010 12:56:40 -0500
- [IGSSTATION-3494]: GANP: data resubission
- Date: Thu, 7 Jan 2010 07:32:29 +0100
- [IGSSTATION-3492]: OPMT: problems with GPS receiver
- Date: Wed, 06 Jan 2010 16:56:24 +0100
- [IGSSTATION-3493]: GRAS station status
- From: Cardaliaguet Jean-Paul
- Date: Wed, 6 Jan 2010 14:21:40 +0100
- [IGSSTATION-3490]: KERG station status
- From: Cardaliaguet Jean-Paul
- Date: Wed, 6 Jan 2010 14:15:09 +0100
- [IGSSTATION-3491]: HARB station status
- From: Cardaliaguet Jean-Paul
- Date: Wed, 6 Jan 2010 14:14:10 +0100
- [IGSSTATION-3489]: THTI station status
- From: Cardaliaguet Jean-Paul
- Date: Wed, 6 Jan 2010 14:02:41 +0100
- [IGSSTATION-3488]: New GNSS stations in the french Caribbean Islands
- Date: Wed, 6 Jan 2010 12:27:36 +0100
- [IGSSTATION-3487]: NANO: data postings resume
- Date: Tue, 5 Jan 2010 13:54:07 -0500
- [IGSSTATION-3486]: BRMU new contact info only
- Date: Tue, 05 Jan 2010 11:32:00 -0500
- [IGSSTATION-3485]: WES2 New Receiver LEICA1200GG Pro
- Date: Tue, 05 Jan 2010 10:20:21 -0500
- [IGSSTATION-3484]: YKRO: data outage
- Date: Tue, 05 Jan 2010 07:57:23 -0700
- [IGSSTATION-3483]: IISC: data outage
- Date: Tue, 05 Jan 2010 07:54:37 -0700
- [IGSSTATION-3482]: SCH2 data flow resumed
- Date: Mon, 4 Jan 2010 14:39:14 -0500
- [IGSSTATION-3481]: BAIE dataflow resumed
- Date: Mon, 4 Jan 2010 11:52:12 -0500
- [IGSSTATION-3480]: NYA1, NYAL: Data gap
- Date: Mon, 4 Jan 2010 15:37:26 +0100
- [IGSSTATION-3479]: YEBE: EXTERNAL H-MASER Frequency problems
- From: Gonzalo López Pedro
- Date: Mon, 4 Jan 2010 13:00:00 +0100
- [IGSSTATION-3478]: LAMA: data missing
- Date: Mon, 04 Jan 2010 12:38:07 +0100
- [IGSSTATION-3477]: HOLB: Data Delivery Delayed
- Date: Sat, 2 Jan 2010 12:42:54 -0500
- [IGSSTATION-3476]: VIS0 resubmitted, DoY 364 and 365
- Date: Fri, 1 Jan 2010 14:52:24 +0100
Mail converted by MHonArc 2.6.10